OpenStudy (anonymous):

Perform the operations indicated. Write your answer in simplest form. 4/5(-15/11)

OpenStudy (anonymous):

have you already tried it?

Parth (parthkohli):

Don't get intimidated by the parentheses. This is just normal multiplication of two fractions, okay?

OpenStudy (anonymous):

yes I just need refresher on how to perform this task

Parth (parthkohli):

\[{4 \over 5} \times {-15 \over 11} \Longrightarrow {4 \times -15 \over 5 \times 11} \]Got it?

OpenStudy (anonymous):

if really don't know how to do this, you'll just multiply numerator to numerator as to deno. to deno..then simplify it...

OpenStudy (anonymous):

ahhh I think so!!!! Thank you so much :)

Parth (parthkohli):

You're welcome! =)

OpenStudy (anonymous):

I need help. Emergency Help on HW: I attended a workshop in Casa Grande Az from 23-27 June. I drove to the workshop from superior to Casa Grande chamber of commerce. Estimate the cost on my transportation for attending the workshop. 1. Explain the strategies you used 2. Give the resources you used to estimate. 3. What information did you need/collect to estimate the solution to the problem. I get graded on the following: One Point to neatly write down the question. Four points for the strategies used and writing them down properly. One point for the correct answer. 2 points for neatness and presentation 2 points is their an algebraic way to solve the problem give your rationale for it?